OpenStudy (anonymous):

solving 3x^2=15x+10 my answer is 15+-sqrt of 105/6. is this correct? i used the quadratic formula. or is there a more efficient way?

OpenStudy (anonymous):

coll... guy ! hmm......

OpenStudy (anonymous):

the 2nd one!

OpenStudy (anonymous):

yup.. having issues..lol

sam (.sam.):

Completing the square will work too, \[3x^2-15x-10=0\] \[3(x^2-5x)-10=0\] \[3(x-\frac{5}{2})^2-\frac{25}{4}(3)-10=0\] \[3(x-\frac{5}{2})^2-\frac{115}{4}=0\] \[3(x-\frac{5}{2})^2=\frac{115}{4}\] \[(x-\frac{5}{2})^2=\frac{115}{12}\] \[x-\frac{5}{2}=\pm\frac{\sqrt{345}}{6}\] \[x=\frac{5}{2}\pm\frac{\sqrt{345}}{6}\]